# GrAviPaSt’s Lens to the Past: Unveiling the Evolution of Filamentary Structures

### v1

@@ 4.2.3 Filament Galaxies Mean Radial Distance
As shown in Figure 5, the mean distance distribution of filament galaxies expands as redshift decreases. At redshift 1, the distribution is more concentrated, but as redshift decreases, it becomes more dispersed, covering a broader range. To further analyze these dynamics, macro-filaments are classified based on the weighted mean velocity of their hosted galaxies. Two categories are defined: macro-filaments with a negative weighted mean velocity, indicating a predominance of matter infall toward the filament skeleton, and macro-filaments with a positive weighted mean velocity, signifying a predominance of outflow motion away from the skeleton. As depicted in Figure 5, the median values of both groups increase with decreasing redshift, and the gap between their medians becomes more distinct over cosmic time. The fraction of macro-filaments exhibiting a negative weighted mean velocity is 0.52, 0.51, and 0.57 for redshifts 0, 0.5, and 1, respectively. The stability of this fraction across redshifts suggests slow dynamical evolution of macro-filaments over cosmic time. Additionally, the mean of weighted mean velocities for the negative group is -153.61 $km/s$, -157.54 $km/s$, and -141.76 $km/s$ at redshifts 0, 0.5, and 1, respectively, whereas for the positive group, it is 175.64 $km/s$, 186.69 $km/s$, and 201.51 $km/s$ across these redshifts. The consistency of mean values within the negative group across different redshifts, along with the gradual decline in mean values for the positive group as redshift decreases, indicates the absorption of matter toward the filament skeletons identified by GrAviPaSt.
A two-sided K-S test was performed to analyze the distribution of mean distance to the skeleton for the negative group. The resulting p-values are $1.73\times 10^{-9}$, $9.37\times 10^{-23}$, and $7.21\times 10^{-12}$ for comparisons between redshifts 0 and 0.5, 0 and 1, and 0.5 and 1, respectively. Similarly, the K-S test applied to the positive group produced p-values of $1.33\times 10^{-4}$, $2.27\times 10^{-19}$, and $4.34\times 10^{-10}$ for the same redshift comparisons. These results demonstrate a statistically significant difference in the distribution of both groups across different redshifts.
@@ 4.2.4 Total Luminosity, Mass and Volume Evolution
From the data in Table 2, it is clear that the total volume of macro-filaments steadily increases from redshift 1 to 0 in both proper and co-moving coordinates. This growth corresponds to the thickening of micro-filaments across these redshifts (see Section 4.1.2). As expected, when accounting for the Hubble flow and the universe’s expansion in the proper volume of macro-filaments, the increase is more pronounced compared to the co-moving coordinate. Furthermore, the total mass and luminosity of filament galaxies demonstrate an overall upward trend, with a significant increase observed between redshifts 1 and 0.5. However, this is followed by a minor decline between redshifts 0.5 and 0.
@@ 5 Summary and Discussion
In addition, the various characteristics of the filaments provide valuable insights into their dynamics. For instance, the velocity patterns of filament galaxies along the length of micro-filaments. Moreover, the mean weighted velocity of these galaxies highlights the overall inflow and outflow of matter relative to the filament skeleton. The observations based on these weighted mean velocities indicates the absorption of matter toward the cosmic filament skeletons (Section 4.2.3). Additionally, the relatively close mean distance of galaxies to the skeleton for both categories of macro-filaments across different redshifts emphasizes the structural consistency of these filamentary structures (Section 4.2.3).
These analyses have been presented to illustrate various aspects of the cosmic filaments and their dynamics. However, further investigation on some of these characteristics is necessary in future studies. The key objectives of this analysis were to assess the effectiveness of the GrAviPaSt method in identifying filaments and to uncover new insights into the evolutionary trends of cosmic filaments from redshift 1 to 0. Nevertheless, it remains challenging to fully assess the success of any filament identification approach due to the lack of a detailed consensus on the definition of cosmic filaments, while different studies have employed diverse methods and criteria to define these structures.

### v2

@@ 4.2.3 Filament Galaxies Mean Radial Distance
As shown in Figure 5, the mean distance distribution of filament galaxies expands as redshift decreases. At redshift 1, the distribution is more concentrated, but as redshift decreases, it becomes more dispersed, covering a broader range. To further analyze these dynamics, macro-filaments are classified based on the weighted mean velocity of their hosted galaxies. Two categories are defined: macro-filaments with a negative weighted mean velocity, indicating a predominance of matter infall toward the filament skeleton, and macro-filaments with a positive weighted mean velocity, signifying a predominance of outflow motion away from the skeleton. As depicted in Figure 5, the median values of both groups increase with decreasing redshift, and the gap between their medians becomes more distinct over cosmic time. The fraction of macro-filaments exhibiting a negative weighted mean velocity is 0.56, 0.49, and 0.5 for redshifts 0, 0.5, and 1, respectively. The stability of this fraction across redshifts suggests slow dynamical evolution of macro-filaments over cosmic time. Additionally, the mean of weighted mean velocities for the negative group is -169.27 $km/s$, -150.09 $km/s$, and -151.06 $km/s$ at redshifts 0, 0.5, and 1, respectively, whereas for the positive group, it is 181.51 $km/s$, 172.73 $km/s$, and 164.37 $km/s$ across these redshifts. The stability of mean values within both groups across different redshifts suggests a slow dynamical evolution of macro-filaments.
Two-sided K–S tests were applied to analyze the distributions of mean distance to the skeleton for both the negative and positive groups, across redshift pairs 0 and 0.5, 0 and 1, and 0.5 and 1. The resulting p-values, as outlined in Table 3, indicate statistically significant differences in the distribution of both groups across different redshifts.
@@ 4.2.4 Total Luminosity, Mass and Volume Evolution
Based on the data in Table 4, the total volume of macro-filaments steadily increases from redshift 1 to 0 in proper coordinates. In co-moving coordinates, the total volume increases from redshift 1 to 0.5 and then decreases from redshift 0.5 to 0. This growth in proper coordinates corresponds to the thickening of micro-filaments across these redshifts (see Section 4.1.2). Additionally, the total mass and luminosity of filament galaxies follow a similar trend to the total co-moving volume, increasing between redshifts 1 and 0.5 and then decreasing between redshifts 0.5 and 0.
@@ 5 Summary and Discussion
In this study we presented GrAviPaSt, a fully deterministic filament-finding algorithm with no free parameters. Its adaptive nature allows it to respond dynamically to the structure and local conditions of the input dataset, producing micro-filaments that reflect the underlying gravitational environment without external tuning. Each phase of the algorithm is directly informed by the data. In the first phase (Section 3.1), Prim’s MST operates solely based on the spatial configuration of the BGG (or BCG) centers. In the second phase (Section 3.2), the A*-like path-finding algorithm depends both on the input dataset, which determines the grid spacing $d$, and on the local geometry of each galaxy group (or cluster) pair, which specifies the coordinate ranges of the corresponding cropped box. Since the cropped box dimensions are determined by the spatial extent of group or cluster members, the gravitational potential computed within each cropped box is unique, and consequently, the traced potential-minimizing path varies from one micro-filament to another. This local responsiveness continues in the third phase (Section 3.3), where the search for filament galaxies is carried out around the skeleton of each micro-filament, with the search radius determined by using the $R_{200}$ values of the associated group (or cluster) pair. The final radial extent of each micro-filament is then determined by the distribution of galaxies along its skeleton, ensuring environment-specific adaptation. Lastly (Section 3.4), the grouping of micro-filaments into macro-filaments proceeds by evaluating the positional, geometric, and gravitational potential characteristics of all micro-filaments connected to each node (group or cluster). This enables the algorithm to adapt to the local environment and robustly identify coherent pairs of micro-filaments to be linked together. Together, these fully data-driven procedures demonstrate the robustness of GrAviPaSt’s results and emphasize the method’s capacity to adapt naturally to diverse cosmic environments.
@@
In addition, the various characteristics of the filaments provide valuable insights into their dynamics. For instance, the velocity patterns of filament galaxies along the length of micro-filaments. Moreover, the mean weighted velocity of these galaxies highlights the overall inflow and outflow of matter relative to the filament skeleton. Additionally, the relatively close mean distance of galaxies to the skeleton for both categories of macro-filaments across different redshifts emphasizes the structural consistency of these filamentary structures (Section 4.2.3).
These analyses aimed to assess the effectiveness of the GrAviPaSt method in identifying filaments and to uncover new insights into the evolutionary trends of cosmic filaments from redshift 1 to 0. Nevertheless, it remains challenging to fully evaluate the success of any filament identification approach due to the absence of a detailed consensus on the definition of cosmic filaments, as different studies have adopted diverse methods and criteria for defining these structures. Also, these analyses have helped illustrate various aspects of cosmic filaments and their dynamics. However, further investigation of these characteristics remains necessary in future studies.
[…]
An alternative and widely used approach to the cosmic web classification, based on the gravitational potential of the dataset, involves analyzing the deformation tensor (i.e., the Hessian of the gravitational potential) and offers a complementary framework to GrAviPaSt. For example, the T-web method (e.g., J. Forero-Romero et al. (2009); E. Ayçoberry et al. (2024)) utilizes the eigenvalues of the deformation tensor to classify the large-scale environment into voids, sheets, filaments, and knots. This technique relies on the local curvature of the potential field to infer structural morphology and identifies filament cells as those with exactly two eigenvalues above a fixed threshold.
While GrAviPaSt also relies on the gravitational potential field, it differs conceptually and operationally from Hessian-based methods. Rather than classifying the cosmic web based on local curvature or eigenvalue thresholds, GrAviPaSt constructs micro-filament skeletons by identifying the path of deepest gravitational potential between galaxy groups. This approach is more trajectory-based and emphasizes the connectivity and continuity of filamentary structures. In this sense, GrAviPaSt complements Hessian-based methods by offering a path-centric, potential-guided reconstruction of filaments, which may be particularly useful for studying their dynamical evolution.
@@
We extend our sincere appreciation to the anonymous referee for their thoughtful evaluation and constructive critique, which have significantly enhanced the clarity, depth, and scholarly rigor of this work.